OpenStudy (anonymous):

√72a^5 simplify the expression

OpenStudy (anonymous):

\[\sqrt{36} * \sqrt{2} * a^2\] \[6\sqrt{2} * a^2\] \[36\sqrt{2} * a^2\]

OpenStudy (anonymous):

thankyou